OpenStudy (anonymous):

To triple both the pressure and the volume of a fixed amount of an ideal gas, you would multiply its absolute temperature by (a) 1 (b) 3 (c) 6 (d) 9

OpenStudy (anonymous):

PV=nRT

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

so I think its 3

OpenStudy (anonymous):

\[\frac{ P*V }{ T }=constant\]

OpenStudy (anonymous):

according to above gas eq, it should be 3*3=9

OpenStudy (anonymous):

answer should be 9

OpenStudy (anonymous):

\[\frac{ P1*V1 }{ T1 }=\frac{ 3P1*3V1 }{ T2 }\]

OpenStudy (anonymous):

\[T2=9T1\]

OpenStudy (anonymous):

@cormacpayne do you think its 9 too

OpenStudy (anonymous):

so you guys are saying it 9 because we have P & V so we will multiply each by 3 that makes it 9

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

I thought it would be 3 since they said you have a fixed amount of gas (meaning the volume is constant), but it would make sense if it was 9 as well.

OpenStudy (anonymous):

i think fixed amount means mass of gas.

OpenStudy (anonymous):

Okay that makes sense. The answer is 9 then.

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

yw